OpenStudy (anonymous):

What is the solution set of |x – 6| = 10?

OpenStudy (anonymous):

-4 and 16

Parth (parthkohli):

Oh, my, goodness!

Parth (parthkohli):

Convert this into a plus/minus equation to start with.\[ x - 6 = \pm 10\]Now, we have two different equations to solve one-by-one: \(x - 6 = 10\) and \(x - 6 = -10\)